39.8g of a mixture of potassium chloride KCl and potassium trioxochlorate(V), KClO3 were heated to a constant mass. If the residue weighed 28.9g, what was the percentage mass of the potassium chloride in the mixture


1
Expert's answer
2022-03-16T09:08:01-0400

KClO3 + 2KCl → 3KClO

"KClO_3"  - x gram

2 moles of Potassium trioxochlorate produces 2 moles of Potassium chloride

"\frac{x}{122.5}" Moles will produce same amount of KCl

"\\""\frac{x}{122.5}\ mole \ KCl=\frac{x}{122.5}\times 74.5\ g\ KCl"

write an equation

"39.8-x+\frac{74.5x}{122.5}=28.9"

x=27.82g

m(KCl)=39.8-27.82=11.98g

"the percentage mass=\frac{11.98}{39.8}\times100=30.10" %